Find parametric equations for the rectangular equation x² + y² - 36 = 0.
x = cos(6t), y = sin(6t), 0 < t < 2π
x = 5 cos(t), y = 5 sin(t), 0 < t < 2π
x = 36 cos(t), y = 36 sin(t), 0 < t < 2π
x = 6 cos(t), y = 6 sin(t), 0 < t < 2π

Answer by ikleyn(52781)

.
Find parametric equations for the rectangular equation x² + y² - 36 = 0.

x = cos(6t), y = sin(6t), 0 < t < 2π
x = 5 cos(t), y = 5 sin(t), 0 < t < 2π
x = 36 cos(t), y = 36 sin(t), 0 < t < 2π
x = 6 cos(t), y = 6 sin(t), 0 < t < 2π       <--- This and only this.